OpenStudy (anonymous):

Solve -6=3x-6y 4x=4+5y

OpenStudy (anonymous):

4x-5y=4 3x-6y=-6

OpenStudy (anonymous):

(6,4)

OpenStudy (anonymous):

oh thanks statlite...i was going so lazy to do the computations:P

OpenStudy (anonymous):

just seeing if i could answer one before armistre or uzma. my mind is shot

OpenStudy (anonymous):

:D thumbs up for u

OpenStudy (anonymous):

i would request admin to install a like button in next upgradation;)

OpenStudy (lone001):

-6=3x-6y -4=-4x+5y multiply first with 3 and second with 4 -24=12x-24y -12=-12x+15y sum side by side -36=-9y y=4 x=6